If two flagpoles are 10m and 70m tall and are 100m apart, find the height of the point where a line from the top of the first to the bottom of the second intersects a line from the bottom of the first to the top of the second.
-Ok i tried setting up the problem in proportions, & drawing it but i have no idea what to do, it is a very confusing question.

Answer by venugopalramana(3286)

If two flagpoles are 10m and 70m tall and are 100m apart, find the height of the point where a line from the top of the first to the bottom of the second intersects a line from the bottom of the first to the top of the second.
LET THE 2 FLAG POLES BE AB AND CD,WITH A AND C REPRESENTING TOPS AND B AND D POINTS ON THE GROUND.LET AD AND BC MEET AT P DRAW PQ PERPENDICULAR TO THE GROUND.WE HAVE
AB=10 FT.
CD=70 FT.
BD=100 FT
WE HAVE TO FIND PQ LET IT BE H.
TRIANGLES DPQ AND DAB ARE SIMILAR.SO
PQ/AB = DQ/DB
H/10=DQ/100
DQ=10H.............I
TRIANGLES BCD AND BPQ ARE SIMILAR .SO
PQ/CD=BQ/BD
H/70=BQ/100
BQ=H*10/7.............II
ADDING EQN.I AND EQN.II
DQ+BQ=10H+10H/7
BD=10H(1+1/7)=80H/7
100=80H/7
H=7*100/80=70/8=8.75 FT.